OpenStudy (anonymous):

how to simplify ((-33Pi)/2) + (1/(1+66Pi)) into -99Pi

OpenStudy (anonymous):

\[\frac{ -33\Pi }{ 2 } + \frac{ 1 }{ 1 + 66\Pi }\]

OpenStudy (anonymous):

the answer is -99Pi

OpenStudy (anonymous):

how do i this to that

OpenStudy (anonymous):

Solve it step by step

OpenStudy (anonymous):

i tried multiplying by the common denominator but it just takes me back to step 1

OpenStudy (anonymous):

\[\frac{ -33\Pi(1+66\Pi) + 2 }{ 2(1+66\Pi) }\]

OpenStudy (anonymous):

If I remember right you would have to multiply everything with pi beside it by 3.14, but thats not a promise. Its been awhile since I've worked with problems including pi

OpenStudy (anonymous):

i c

OpenStudy (anonymous):

btw the original equation was \[y = \sin \left[ x - \tan(\frac{ \Pi }{ 4 }x ^{66}) \right] + x ^{\frac{ 1 }{ 1+66\Pi }}\]

OpenStudy (anonymous):

with 1 plugged in

OpenStudy (anonymous):

i had to takes its derivative and then narrow it down to what i have now

OpenStudy (anonymous):

the answer is -99Pi, but i need to know how it simplified to that

OpenStudy (anonymous):

\[y'(1) = \cos \left[ x- \tan(\frac{ \Pi }{ 4 }x ^{66}) \right](1 - \sec ^{2}(\frac{ \Pi }{ 4 }x ^{66}))(\frac{ 33\Pi }{ 2 }x ^{65}) + \frac{ 1 }{ 1 + 66\Pi }x ^{\frac{ 1 }{ x+66\Pi }-1}\]

OpenStudy (anonymous):

the x at the end is not an exponent

OpenStudy (anonymous):

if you cant see the exponent at the end, it says\[\frac{ 1 }{ 1+66\Pi }\]

OpenStudy (anonymous):

subtraced by 1

OpenStudy (anonymous):

actually i may have taken the derivative wrong

zepdrix (zepdrix):

woahh nasty looking problem XD

OpenStudy (anonymous):

yea